# Convergence and Optimality of the EM Algorithm Under Multi-Component Gaussian Mixture Models

### v1

@@ 1.2 Our contributions
We now summarize our main results which include analyses of the EM algorithm at both the population level and the sample level. For the former, we establish a contraction region for EM, under what is, to our best knowledge, the mildest separation condition known for $L\geq 3$. For the latter, our analysis is based on two complementary approaches, reflecting different perspectives on EM: parameter estimation and clustering. The first approach is more suitable when the separation is small to moderate, while the second is particularly effective when the separation is moderate to large. To distinguish between the true model parameters and their iterative estimates, we denote the true parameter by $\theta^{*}=(\pi^{*},M^{*},\Sigma^{*})$.
@@ 1.2.1 Convergence of the population level EM algorithm
the population level EM iterates $\theta^{(t)}=(\pi^{(t)},M^{(t)},\Sigma^{(t)})$ with $t\geq 0$, once initialized within a neighborhood of $\theta^{*}$, converges as: for all $t\geq 1$ and some contraction rate $\kappa\in(0,1)$,
@@
The distances between the parameters in $\theta^{(t)}$ and $\theta^{*}$ are given in Eqs. 3.2 and 3.3. The convergence in (1.4) is at least linear as $t\to\infty$ and becomes superlinear once $\Delta_{\min}$ increases, since our results show that $\kappa$ decays exponentially in $\Delta_{\min}$ (see, (3.6)). Regarding the initialization, Theorem 2 shows that (1.4) requires
@@
for some absolute constants $0\leq 2c_{\mu}+\sqrt{2}c_{\Sigma}<\sqrt{2}-1$. For known $\Sigma^{*}$, we have $\Sigma^{(t)}=\Sigma^{*}$ for all $t\geq 0$ in (1.4) and $c_{\Sigma}=0$ in (1.5). Moreover, our 3 states that the contraction rate $\kappa$ for known $\Sigma^{*}$ becomes faster by a factor of $\Delta_{\max}$. For unknown $\Sigma^{*}$, we note in Section 2.1 that the EM iterates $\Sigma^{(t)}$ depend on $\mathbb{E}_{\theta^{*}}[XX^{\top}]$, $\pi^{(t)}$ and $M^{(t)}$, which suggests a simple way to choose $\Sigma^{(0)}$. For such a choice, 4 of Section 3 shows that $d(\Sigma^{(0)},\Sigma^{*})\leq c_{\Sigma}$ in (1.5) is satisfied, provided that the requirements on $d(\pi^{(0)},\pi^{*})$ and $d(M^{(0)},M^{*})$ in (1.5) are tightened by a multiplicative factor of $1/\Delta_{\max}$ on the right-hand side.
To the best of our knowledge, this is the first convergence result of the population level EM algorithm under the $L$-GMM with $L\geq 3$, unknown $\pi^{*}$, and either known or unknown $\Sigma^{*}$. As mentioned earlier, for known $\pi^{*}$ and $\Sigma^{*}=\mathbf{I}_{d}$, [50, 51] analyze the population level EM and its variant under (1.2), a separation condition that is strictly stronger than ours in (1.3), often substantially so when $L$ and $d$ are moderate to large. Their initialization requirement on $M^{(0)}$ is similar to ours in (1.5), although derived under the stronger separation condition in addition to known $\pi^{*}$ and $\Sigma^{*}=\mathbf{I}_{d}$.
@@
Throughout this paper, we use $\bm{\pi}$, $\mathbf{M}$, and $\bm{\Sigma}$ to denote operators, in order to distinguish them from their vector or matrix counterparts $\pi$, $M$, and $\Sigma$. By further using the self-consistency of EM in Section C.1, that is, $\theta^{*}=(\bm{\pi}(\theta^{*}),\mathbf{M}(\theta^{*}),\bm{\Sigma}(\theta^{*}))$, bounding $d(\pi^{(t+1)},\pi^{*})$ boils down to controlling $d(\bm{\pi}(\theta^{(t)}),\bm{\pi}(\theta^{*}))$. This, in turn, amounts to establishing the Lipschitz continuity of the operator $\theta\mapsto\bm{\pi}(\theta)$, and similarly for $\theta\mapsto\mathbf{M}(\theta)$ and $\theta\mapsto\bm{\Sigma}(\theta)$, which constitutes the main challenge in our analysis. Our proof first re-parametrizes $\theta$ by $\omega(\theta)=(\pi,M,J(\theta))$ with $J(\theta):=\Sigma^{-1}M$ as such reparametrization alleviates the difficulty of direct manipulation of $\Sigma^{-1}$. Then in Theorem 24 of Section D.2 we establish the Lipschitz continuity of $\omega(\theta)\mapsto\bm{\pi}(\omega(\theta))$ and $\omega(\theta)\mapsto\mathbf{M}(\omega(\theta))$, by quantifying the partial derivatives of $\bm{\pi}(\cdot)$ and $\mathbf{M}(\cdot)$ with respect to $\pi$, $M$, and $J(\theta)$ uniformly for all $\theta$ satisfying (1.5) in lieu of $\theta^{(0)}$. Finally, combined with the Lipschitz property of $\theta\mapsto J(\theta)$ established in Lemma 18, the conclusion in (1.4) follows. Allowing $L\geq 3$, unknown $\pi^{*}$ and $\Sigma^{*}$, and avoiding strong separation conditions significantly complicates the analysis. We refer the reader to the end of Section 3 for a more detailed technical discussion.
@@ 1.2.2 Convergence of the sample level EM algorithm
We emphasize that this is a different decomposition from the one used in the literature, such as [4, 50, 9], and is crucial to obtain minimax optimal rates. We refer to the beginning of Section 4.1 for detailed explanations.
In 7 of Section 4.1, we control the first term on the right-hand side by establishing the Lipschitz properties of the sample level operators $\theta\mapsto\widehat{\bm{\pi}}(\theta)$, $\theta\mapsto\widehat{\mathbf{M}}(\theta)$, and $\theta\mapsto\bm{\widehat{\Sigma}}(\theta)$, which readily yield: with probability $1-\mathcal{O}(n^{-1})$, for all $\theta$ satisfying (1.5) in lieu of $\theta^{(0)}$,
@@
It is for this reason, this analysis is only suitable for small to moderate separation. We pause here to clarify that although our analysis significantly relaxes the upper bound restriction on $\Delta_{\max}$ compared to existing results, the condition imposed in (1.7) remains inevitable, as it is rooted in directly proving the convergence of $\widehat{\theta}^{(t)}$ to $\theta^{*}$, even in the case $L=2$. Nevertheless, as we explain below, this dependence can be avoided by adopting a different analytical approach.
Regarding the second term $d(\widehat{\bm{h}}(\theta^{*}),\bm{h}(\theta^{*}))$ in (1.6), we derive in Theorem 5 of Section 4.1 concentration inequalities between the operators $\widehat{\bm{\pi}}$, $\widehat{\mathbf{M}}$, and $\bm{\widehat{\Sigma}}$ and their population-level counterparts, evaluated at $\theta^{*}$. To obtain deviation bounds with near-optimal dependence on $\Delta_{\max}$, $\Delta_{\min}$, and $\pi_{\min}^{*}$, we conduct a separate analysis for the regimes $\Delta_{\min}\gtrsim\log n$ and $\Delta_{\min}\lesssim\log n$, as standard concentration inequalities alone are insufficient for capturing the sharp dependence on these parameters.
Finally, by combining the bounds for both terms in (1.6), we establish in Theorem 9 of Section 4.1 that the EM iterates $\widehat{\theta}^{(t)}$, initialized with $\widehat{\theta}^{(0)}$ satisfying (1.5) in lieu of $\theta^{(0)}$, satisfy: with high probability, for all $t\geq C\log n$,
@@
The main idea of our proof consists of two steps: in the first step we introduce a surrogate loss related to the community detection problem under model (1.1) and establish its convergence; in the second step we translate this convergence into a convergence rate for estimating $\theta^{*}$. The surrogate loss, $\phi(\theta)$ given by (4.10), characterizes the discrepancy between the posterior probabilities $\mathbb{P}_{\theta}(Y_{i}=\ell\mid X_{i})$ in the E-step at parameter value $\theta$, and the true labels $1\{Y_{i}=\ell\}$ for all $i\in[n]$ and $\ell\in[L]$. It also captures the degree of separation between the Gaussian components, thereby reflecting the benefit of increased separation. In Theorems 11 and 10 of Section 4.2, we establish the convergence of $\phi(\widehat{\theta}^{(t)})$ along the EM iterates $\widehat{\theta}^{(t)}$: with probability at least $1-4n^{-1}-\exp(-c\Delta_{\min})$ for some absolute constant $c>0$,
@@
for both known and unknown $\Sigma^{*}$. In addition to the initialization condition in (1.5), the above result admits an alternative requirement on $\widehat{\theta}^{(0)}$ in terms of $\phi(\widehat{\theta}^{(0)})$, which can be linked to preliminary estimates of the labels. On the other hand, unlike the contraction-based argument in our first approach, the convergence result in (1.9) requires only a lower bound on $\Delta_{\min}$ and, in fact, improves with larger values of $\Delta_{\min}$; see, details in Theorems 11 and 10.
@@
Finally, by combining the results from both approaches, we have established that the EM iterates satisfy the bounds in (1.8) after $\mathcal{O}(\log n)$ iterations, and is also minimax optimal up to logarithmic factors, for all $\Delta_{\min}$ satisfying (1.3). Although our results are established for $L\geq 3$, it is worth noting that they also improve upon existing results in the case of $L=2$ with unknown $\pi^{*}$ and $\Sigma^{*}$; see, 15 and its subsequent discussion.
@@ 3 Convergence of the population level EM algorithm
The smallest separation, $\Delta_{\min}$, is known as a key quantity of learning Gaussian location mixtures. Regarding $\Delta_{\max}$, we have $\Delta_{\max}=\Delta_{\min}$ when $L=2$. For $L\geq 3$, [25] finds that local search procedures may fail to find the global optimum of the likelihood function when one mixture component is significantly farther from the others. This suggests that $\Delta_{\max}$, or more precisely the ratio $\Delta_{\max}/\Delta_{\min}$, influences the difficulty of learning Gaussian mixtures. Another well-known quantity that affects the learning is the smallest mixing probabilities $\pi_{\min}^{*}=\min_{\ell\in[L]}\pi_{\ell}^{*}.$ Our results below characterize the effect of $\Delta_{\min}$, $\Delta_{\max}/\Delta_{\min}$ and $1/\pi_{\min}^{*}$ on the convergence of the EM algorithm. Finally, since we can always subtract the marginal mean of $X$, we assume $\mathbb{E}[X]=\sum_{\ell=1}^{L}\pi_{\ell}^{*}\mu_{\ell}^{*}=0$ throughout our analysis.
@@
Condition (3.4) states the required separation condition between Gaussian components for the population level EM iterates to converge. It characterizes the effect of both the smallest mixing probability $\pi_{\min}^{*}$ and the ratio between the largest and smallest separations among Gaussian components, $\Delta_{\max}/\Delta_{\min}$. We emphasize that such dependence is particularly informative for $L\geq 3$. In the case of $L=2$, we have $\Delta_{\max}=\Delta_{\min}$, and it is standard to assume that $\pi_{1},\pi_{2}\in(c,1-c)$ for some $c\in(0,1/2)$, under which condition (3.4) reduces to $\Delta_{\min}\geq C$. This is assumed in [4] for symmetric, isotropic setting, and in [9] when $\pi^{*}$ and $\Sigma^{*}$ are unknown. For $L\geq 3$, even when $\Delta_{\max}\asymp\Delta_{\min}$ and $\pi_{\min}^{*}\asymp 1/L$, condition (3.4) simplifies to $\Delta_{\min}\geq C\log(L)$, still reflecting the effect of $L$ on the required separation. To the best of our knowledge, condition (3.4) is the weakest separation requirement by far for analyzing the convergence of EM under the $L$-GMM with $L\geq 3$. We refer to Section 4.4 for comparison with the separation conditions of clustering-based algorithms.
@@
The requirement in (3.5) on $M^{(0)}$ becomes milder as the separation $\Delta_{\min}$ increases. This is consistent with existing results [4] for $L=2$, and [50, 51] for $L\geq 3$. Compared to the latter, their initialization requirement is comparable to ours, although derived under known $\pi^{*}$, $\Sigma^{*}=\mathbf{I}_{d}$ and a stronger separation condition. Since we allow both $\pi^{*}$ and $\Sigma^{*}$ unknown, our initialization requirement in (3.5) also depends on $\pi^{(0)}$ and $\Sigma^{(0)}$. The number $1/2$ in $d(\pi^{(0)},\pi)$ can be replaced by any nonnegative value less than $1$. As shown in Table 1, when $\pi^{*}$ is unknown, no prior work establishes the convergence of the EM algorithm for the $L$-GMM with $L\geq 3$ even for known $\Sigma^{*}$. Later in Section 4.2 we introduce an alternative initialization scheme that also enjoys provable guarantees.
@@
From the technical perspective, establishing the Lipschitz continuity in (3.10) for both $\bm{\pi}_{\ell}(\omega(\theta))$ and $\bm{\mu}_{\ell}(\omega(\theta))$ is the most challenging part. It requires quantifying the partial derivatives of $\gamma_{\ell}(x;\omega(\theta))$ with respect to $\pi$, $M$, and $J(\theta)$ uniformly over a neighborhood of $\omega(\theta^{*})$. The proofs of (3.10) are provided in Section C.3.1 and Section C.3.2, along with the supporting technical lemmas in Lemmas 21, 23 and 22 of Section C.4.

### v2

@@ 1.2 Our contributions
We now summarize our main results which include analyses of the EM algorithm at both the population level and the sample level. For the former, we establish a contraction region for EM while for the latter, our analysis is based on two complementary approaches, reflecting different perspectives on EM: parameter estimation and clustering. The first approach is more suitable when the separation is small to moderate, while the second is particularly effective when the separation is moderate to large. To distinguish between the true model parameters and their iterative estimates, we denote the true parameter by $\theta^{*}=(\pi^{*},M^{*},\Sigma^{*})$.
@@ 1.2.1 Convergence of the population level EM algorithm
the population level EM iterates $\theta^{(t)}=(\pi^{(t)},M^{(t)},\Sigma^{(t)})$ with $t\geq 0$, once initialized within a neighborhood of $\theta^{*}$, converges: for all $t\geq 1$ and some contraction rate $\kappa\in(0,1)$,
@@
The distances between the parameters in $\theta^{(t)}$ and $\theta^{*}$ are given in Eqs. 3.2 and 3.3. The convergence in (1.3) is at least linear as $t\to\infty$ and becomes superlinear once $\Delta_{\min}$ increases, since our results show that $\kappa$ decays exponentially in $\Delta_{\min}$ (see, (3.6)). To the best of our knowledge, this is the first convergence result of the population level EM algorithm under the $L$-GMM with $L\geq 3$, unknown $\pi^{*}$ and unknown $\Sigma^{*}$. Compared to the required separation when $\Sigma^{*}=\mathbf{I}_{d}$, we only pay the price of a logarithmic factor in $\Delta_{\max}$ for estimating $\Sigma^{*}$. Regarding the initialization, Theorem 2 shows that (1.3) requires
@@
for some absolute constants $0\leq 2c_{\mu}+\sqrt{2}c_{\Sigma}<\sqrt{2}-1$. The existing initialization requirements on $M^{(0)}$ are similar to ours in (1.4), although derived under known $\pi^{*}$ and $\Sigma^{*}=\mathbf{I}_{d}$. For known $\Sigma^{*}$, we have $\Sigma^{(t)}=\Sigma^{*}$ for all $t\geq 0$ in (1.3) and $c_{\Sigma}=0$ in (1.4). Moreover, our 3 states that the contraction rate $\kappa$ for known $\Sigma^{*}$ becomes faster by a factor of $\Delta_{\max}$. For unknown $\Sigma^{*}$, we note in Section 2.1 that the EM iterates $\Sigma^{(t)}$ depend on $\mathbb{E}_{\theta^{*}}[XX^{\top}]$, $\pi^{(t)}$ and $M^{(t)}$, which suggests a simple way to choose $\Sigma^{(0)}$. For such a choice, 4 of Section 3 shows that $d(\Sigma^{(0)},\Sigma^{*})\leq c_{\Sigma}$ in (1.4) is satisfied, provided that the requirements on $d(\pi^{(0)},\pi^{*})$ and $d(M^{(0)},M^{*})$ in (1.4) are tightened by a multiplicative factor of $1/\Delta_{\max}$ on the right-hand side.
To the best of our knowledge, this is the first convergence result of the population level EM algorithm under the $L$-GMM with $L\geq 3$, unknown $\pi^{*}$, and either known or unknown $\Sigma^{*}$. As mentioned earlier, for known $\pi^{*}$ and $\Sigma^{*}=\mathbf{I}_{d}$, [yan2017convergence, Zhao2020] analyze the population level EM and its variant under a separation condition that is strictly stronger than ours in (1.2), often substantially so when $L$ and $d$ are moderate to large. Their initialization requirement on $M^{(0)}$ is similar to ours in (1.4), although derived under the stronger separation condition in addition to known $\pi^{*}$ and $\Sigma^{*}=\mathbf{I}_{d}$.
@@
Throughout this paper, we use $\bm{\pi}$, $\mathbf{M}$, and $\bm{\Sigma}$ to denote operators, in order to distinguish them from their vector or matrix counterparts $\pi$, $M$, and $\Sigma$. By further using the self-consistency of EM in Section C.1, that is, $\theta^{*}=(\bm{\pi}(\theta^{*}),\mathbf{M}(\theta^{*}),\bm{\Sigma}(\theta^{*}))$, bounding $d(\pi^{(t+1)},\pi^{*})$ boils down to controlling $d(\bm{\pi}(\theta^{(t)}),\bm{\pi}(\theta^{*}))$. This, in turn, amounts to establishing the Lipschitz continuity of the operator $\theta\mapsto\bm{\pi}(\theta)$, and similarly for $\theta\mapsto\mathbf{M}(\theta)$ and $\theta\mapsto\bm{\Sigma}(\theta)$, which constitutes the main challenge in our analysis. Our proof first re-parametrizes $\theta$ by $\omega(\theta)=(\pi,M,J(\theta))$ with $J(\theta):=\Sigma^{-1}M$ as such reparametrization alleviates the difficulty of direct manipulation of $\Sigma^{-1}$. Then in Theorem 23 of Section D.2 we establish the Lipschitz continuity of $\omega(\theta)\mapsto\bm{\pi}(\omega(\theta))$ and $\omega(\theta)\mapsto\mathbf{M}(\omega(\theta))$, by quantifying the partial derivatives of $\bm{\pi}(\cdot)$ and $\mathbf{M}(\cdot)$ with respect to $\pi$, $M$, and $J(\theta)$ uniformly for all $\theta$ satisfying (1.4) in lieu of $\theta^{(0)}$. Finally, combined with the Lipschitz property of $\theta\mapsto J(\theta)$ established in Lemma 17, the conclusion in (1.3) follows. Allowing $L\geq 3$, unknown $\pi^{*}$ and $\Sigma^{*}$, and avoiding strong separation conditions significantly complicates the analysis. We refer to the end of Section 3 for a more detailed technical discussion.
@@ 1.2.2 Convergence of the sample level EM algorithm
We emphasize that this is a different decomposition from the one used in the literature, such as [EM2017, yan2017convergence, cai2019chime], and is crucial to obtain minimax optimal rates. We refer to the beginning of Section 4.1 for detailed explanations.
In 6 of Section 4.1, we control the first term on the right-hand side by establishing the Lipschitz properties of the sample level operators $\theta\mapsto\widehat{\bm{\pi}}(\theta)$, $\theta\mapsto\widehat{\mathbf{M}}(\theta)$, and $\theta\mapsto\bm{\widehat{\Sigma}}(\theta)$, which readily yield: with probability $1-\mathcal{O}(n^{-1})$, for all $\theta$ satisfying (1.4) in lieu of $\theta^{(0)}$,
@@
It is for this reason, this analysis is only suitable for small to moderate separation. We pause here to clarify that although our analysis significantly relaxes the upper bound restriction on $\Delta_{\max}$ compared to existing results [yan2017convergence, Zhao2020], the condition imposed in (1.6) remains inevitable, as it is rooted in directly proving the convergence of $\widehat{\theta}^{(t)}$ to $\theta^{*}$. Nevertheless, as we explain below, this dependence can be avoided by adopting a different analytical approach.
Regarding the second term $d(\widehat{\bm{h}}(\theta^{*}),\bm{h}(\theta^{*}))$ in (1.5), we derive in Theorem 5 of Section 4.1 concentration inequalities between the operators $\widehat{\bm{\pi}}$, $\widehat{\mathbf{M}}$, and $\bm{\widehat{\Sigma}}$ and their population-level counterparts, evaluated at $\theta^{*}$. By combining the bounds for both terms in (1.5), we establish in Theorem 8 of Section 4.1 that the EM iterates $\widehat{\theta}^{(t)}$, initialized with $\widehat{\theta}^{(0)}$ satisfying (1.4) in lieu of $\theta^{(0)}$, satisfy: with high probability, for all $t\geq C\log n$,
@@
The main idea of our proof consists of two steps: in the first step we introduce a surrogate loss related to the community detection problem under model (1.1) and establish its convergence; in the second step we translate this convergence into a convergence rate for estimating $\theta^{*}$. The surrogate loss, $\phi(\theta)$ given by (4.9), characterizes the discrepancy between the posterior probabilities $\mathbb{P}_{\theta}(Y_{i}=\ell\mid X_{i})$ in the E-step at parameter value $\theta$, and the true labels $1\{Y_{i}=\ell\}$ for all $i\in[n]$ and $\ell\in[L]$. It also captures the degree of separation between the Gaussian components, thereby reflecting the benefit of increased separation. In Theorems 10 and 9 of Section 4.2, we establish the convergence of $\phi(\widehat{\theta}^{(t)})$ along the EM iterates $\widehat{\theta}^{(t)}$: with probability at least $1-4n^{-1}-\exp(-c\Delta_{\min})$ for some absolute constant $c>0$,
@@
for both known and unknown $\Sigma^{*}$. In addition to the initialization condition in (1.4), the above result admits an alternative requirement on $\widehat{\theta}^{(0)}$ in terms of $\phi(\widehat{\theta}^{(0)})$, which can be linked to preliminary estimates of the labels. On the other hand, unlike the contraction-based argument in our first approach, the convergence result in (1.8) requires only a lower bound on $\Delta_{\min}$ and, in fact, improves with larger values of $\Delta_{\min}$; see, details in Theorems 10 and 9.
@@
Finally, by combining the results from both approaches, we have established that the EM iterates satisfy the bounds in (1.7) after $\mathcal{O}(\log n)$ iterations, and is also minimax optimal up to logarithmic factors, for all $\Delta_{\min}$ satisfying (1.2). Although our results are established for $L\geq 3$, it is worth noting that they also improve upon existing results in the case of $L=2$ with unknown $\pi^{*}$ and $\Sigma^{*}$; see, 14 and its subsequent discussion.
@@ 3 Convergence of the population level EM algorithm
The smallest separation, $\Delta_{\min}$, is known as a key quantity of learning Gaussian location mixtures. Regarding $\Delta_{\max}$, we have $\Delta_{\max}=\Delta_{\min}$ when $L=2$. For $L\geq 3$, [jin2016local] finds that local search procedures may fail to find the global optimum of the likelihood function when one mixture component is significantly farther from the others. This suggests that $\Delta_{\max}$, or more precisely the ratio $\Delta_{\max}/\Delta_{\min}$, influences the difficulty of learning Gaussian mixtures. Another well-known quantity that affects the learning is the smallest mixing probabilities $\pi_{\min}^{*}=\min_{\ell\in[L]}\pi_{\ell}^{*}.$ Our results below characterize the effect of $\Delta_{\min}$, $\Delta_{\max}/\Delta_{\min}$ and $1/\pi_{\min}^{*}$ on the convergence of the EM algorithm. Finally, since we can always subtract the marginal mean of $X$, we assume $\mathbb{E}[X]=\sum_{\ell=1}^{L}\pi_{\ell}^{*}\mu_{\ell}^{*}=0$ throughout our analysis.
@@
Condition (3.4) states the required separation condition between Gaussian components for the population level EM iterates to converge. It characterizes the effect of both the smallest mixing probability $\pi_{\min}^{*}$ and the ratio between the largest and smallest separations among Gaussian components, $\Delta_{\max}/\Delta_{\min}$. We emphasize that such dependence is particularly informative for $L\geq 3$. In the case of $L=2$, we have $\Delta_{\max}=\Delta_{\min}$, and it is standard to assume that $\pi_{1},\pi_{2}\in(c,1-c)$ for some $c\in(0,1/2)$, under which condition (3.4) reduces to $\Delta_{\min}\geq C$. This is assumed in [EM2017] for symmetric, isotropic setting, and in [cai2019chime] when $\pi^{*}$ and $\Sigma^{*}$ are unknown. For $L\geq 3$, even when $\Delta_{\max}\asymp\Delta_{\min}$ and $\pi_{\min}^{*}\asymp 1/L$, condition (3.4) simplifies to $\Delta_{\min}\geq C\log(L)$, still reflecting the effect of $L$ on the required separation. Such logarithmic dependence on $L$ is known to be necessary for achieving polynomial sample complexity guarantees in the isotropic setting [regev2017learning]. Our condition in (3.4) further shows that when $\Sigma^{*}$ is unknown, the additional price to pay is only a logarithmic factor in $\Delta_{\max}$. We refer to Section 4.4 for comparison with the separation conditions of clustering-based algorithms.
@@
The requirement in (3.5) on $M^{(0)}$ becomes milder as the separation $\Delta_{\min}$ increases. This is consistent with existing results [EM2017] for $L=2$, and [yan2017convergence, Zhao2020, segol2021improved, kwon2020algorithm] for $L\geq 3$. Compared to the latter, their initialization requirement is comparable to ours, although derived under known / unknown $\pi^{*}$ and $\Sigma^{*}=\mathbf{I}_{d}$. Since we allow both $\pi^{*}$ and $\Sigma^{*}$ unknown, our initialization requirement in (3.5) also depends on $\pi^{(0)}$ and $\Sigma^{(0)}$. The number $1/2$ in $d(\pi^{(0)},\pi)$ can be replaced by any nonnegative value less than $1$. Later in Section 4.2 we introduce an alternative initialization scheme that also enjoys provable guarantees. Detailed discussion on practical choices of initialization is given in Section 5.1.
@@
From the technical perspective, establishing the Lipschitz continuity in (3.10) for both $\bm{\pi}_{\ell}(\omega(\theta))$ and $\bm{\mu}_{\ell}(\omega(\theta))$ is the most challenging part. It requires quantifying the partial derivatives of $\gamma_{\ell}(x;\omega(\theta))$ with respect to $\pi$, $M$, and $J(\theta)$ uniformly over a neighborhood of $\omega(\theta^{*})$. The proofs of (3.10) are provided in Section C.3.1 and Section C.3.2, along with the supporting technical lemmas in Lemmas 20, 22 and 21 of Section C.4.